OpenStudy (anonymous):

I am having a whole lot of problems with this Math Lab I'm doing. Help please. 4x+5y=-13 2y+13=x use substitution to solve the system

OpenStudy (anonymous):

what are you trying to solve for

OpenStudy (anonymous):

both x and y

OpenStudy (anonymous):

kay

OpenStudy (anonymous):

2y + 13 =x 2y-x=-13 2y-x=4x + 5y 5x=-3y

OpenStudy (anonymous):

do you still need help

OpenStudy (anonymous):

could you do the rest yourself

OpenStudy (anonymous):

give me a minute plese

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

trying to c if I was doing it right

OpenStudy (anonymous):

I should be ok now I guess. Thank u.

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

y=-1

OpenStudy (anonymous):

that answer was incorrect

OpenStudy (anonymous):

Still working on it